# Super-Tonks-Girardeau quench of dipolar bosons in a one-dimensional optical lattice

### v1

@@ VI Conclusions
We have investigated the quench dynamics of strongly-interacting bosons in an optical lattice for a wide range of dipolar interaction strengths and different fillings, generalizing the STG quench procedure to the presence of long-range interactions and more exotic initial states. Our protocol consisted in preparing an initial state of bosons with strong contact repulsions and dipolar attractions, and then quenching the strength of the contact interactions to the strongly negative (attractive) regime. We have calculated numerous observables to study the time evolution, including one-body density, reduced one-body and two-body densities, entropy measures, and natural occupations. By analyzing these quantities, we have mapped out the effect of different dipolar interaction strengths on the stability of the post-quench states.
As expected, starting from a unit-filled Mott state at zero dipolar interactions we recollect the robustness of the original STG quench, where the localized Mott peaks remain unperturbed for a very long time. At weak dipolar interactions, this stability is retained only for a shorter time. By increasing the dipolar interactions further, the initial bosonic configuration condenses into a cluster state. The cluster state is much more sensitive to the quench procedure and rapidly evaporates into a chaotic state well-described by a thermal Gaussian orthogonal ensemble that maximizes entropy measures. A similar fate is observed by starting from doubly filled states, be them in dimerized Mott configurations (low dipolar interactions) or cluster states (high dipolar interactions). The state rapidly evaporates ejecting particles with complex many-body dynamics that leads to complete fragmentation and thermalization. By examining measure of correlations, we find that the cause for this behavior lies within the structure of the initial states, that exhibit pockets of correlations within the dimers or within different peaks of the cluster states. These pockets of correlations spread upon quenching and lead to instability and eventual evaporation.
However, we have also established that the pathway towards thermalization can be drastically altered by tuning the magnitude of the interactions and of the optical lattice depth. It is possible to obtain prethermal states that survive for longer times before the eventual entropy saturation. Empirically, we find that this behavior is obtained using moderate dipolar interactions strengths in the order of up to one recoil energy. This finding is consistent with previous literature predictions for the appearance of prethermalization due to long-range interactions.
@@
Our study sheds light onto the behavior and stability of strongly correlated bosonic states in the presence of long-ranged interactions. It also showcases the potential of harnessing such long-range interactions and quenching protocols to achieve novel pathways to the creation and control of excited quantum states of matter. These questions are particularly timely since the systems we have investigated should be easily engineered in state-of-the-art ultracold experiments, e.g. with magnetic quantum gases [72]. The complex atomic structure and the large dipole moment of these systems facilitate the control of the interatomic interaction strength, specially the relative strength of contact over dipolar interactions. This will lead to further exploration of few and many-body physics.
@@
In the future, we envision that similar approaches could be employed to study the complex interplay of different energy scales introduced by the competition between localizing optical lattices, local (contact) interactions, and long-range interactions, which could lead to a multitude of exotic quantum states. Our study could also be extended to analyze the stability of other fermionization and crystallization phenomena, potentially in higher dimensions or fostered by other kind of long-range interactions, e.g. mediated by cavities [73, 74]. This last setup is particularly promising as it has been shown to give rise to a plethora of intriguing equilibrium and out-of-equilibrium phases of matter, such as higher-order superfluidity, Mott phases, and limit cycles [75, 76, 77, 78]. It would be interesting to analyze whether the interplay of such phases with dipolar interactions could lead to a stabilization of prethermal states. Another interesting route to pursue is the connection with the accuracy of dipolar quantum simulators [79]. Our method solves the many-body Schrödinger equation in the continuum, and thus should be an accurate representation of the physics at play in the actual quantum simulators and not just in the simulated lattice models (e.g. extended Bose-Hubbard models) studied elsewhere in the literature. For example, a relevant question to ask is to what extend a similar STG quench would lead the physics to deviate from that of a lattice model as a function of dipolar interaction strength, optical lattice depth, and other system parameters. It could also include the investigation of quantum many-body scar states which can give rise to infinitely long-lived coherent dynamics under quantum quenches from well controlled initial states [80]. Another open avenue for future studies is the impact of long-range interactions on quantum holonomy – where parameters are tuned to realized a cycle starting from the free system, to Tonks-Girardeau and super-Tonks-Girardeau regimes, and then back to the free system [81]. The comparison between sudden quench and adiabatic quench will help to understand the underlying correlation dynamics. Furthermore, continuous quenches will provide information on defect dynamics and criticality by means of Kibble-Zurek formalism [82, 83].
@@ Appendix A MCTDH-X
In this appendix, we briefly review the numerical method used to obtain the time evolution of the few-boson systems presented in the main text. We employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles implemented by the MCTDH-X software [49, 50, 51, 53, 52, 54].
@@ Appendix B System parameters
[…]
@@ B.2 Energies
The unit of energy $\bar{E}$ is effectively defined in terms of the recoil energy of the optical lattice, i.e. $E_{r}\equiv\frac{\hbar^{2}k_{0}^{2}}{2m}\approx 2.204\times 10^{-30}$ J with $m\approx 51.941$ Da the mass of ${}^{52}$r atoms, a dipolar species amply used in experiments [88, 89]. More specifically, we define the unit of energy as $\bar{E}\equiv\frac{\hbar^{2}}{m\bar{L}^{2}}=\frac{2E_{r}}{\pi^{2}}\approx 4.466\times 10^{-31}$ J.
In typical experiments, the optical lattice depth is varied in regimes of up to several tens of recoil energies (in particular when lattice physics is quantum simulated). In our simulations, we probe similar regimes of $V_{0}=40\bar{E}\approx 8E_{r}$ or $V_{0}=100\bar{E}\approx 20E_{r}$. The on-site interactions are kept at fixed magnitude $|g_{0}|=400\bar{E}\approx 81E_{r}$, unless otherwise state. During the quench, we change only their sign (from repulsive to attractive). When we turn on dipolar interactions, we probe regimes up to strength $|g_{d}|=45\bar{E}\approx 9E_{r}$, although in the main text we reported only results up to $|g_{d}|=10\bar{E}\approx 2E_{r}$ since the physics does not change qualitatively with stronger dipolar interactions.
@@ B.3 Time
The unit of time is also defined from the unit of length, as $\bar{t}\equiv\frac{m\hat{L}^{2}}{\hbar}=\frac{m\lambda_{0}^{2}}{\hbar}=0.5903\times 10^{-4}$ s $=59.03$ $\mu$s. In our simulations, we ran time evolutions up until around $t\approx 10\bar{t}\approx 590.25$ ms. However, in thermalizing phases, the onset of thermalization is very rapid. Thus, we typically restrict our ranges to show only results up to $t\approx 1\bar{t}\approx 59.03$ $\mu$s.
@@
potential depth $V=400.0\bar{E}\approx 81\>E_{r}$
on-site repulsion $|g_{0}|=400.0\bar{E}\approx 81\>E_{r}$
dipolar interaction e.g. $g_{d}=-1.0\bar{E}=-0.203\>E_{r}$
@@
Appendix C Initial states
In this section we show the density profile for the initial states used in the quench protocol and studied in the main text. The initial states for the unit filling case can be seen in Fig. 12 while the ones for the double filling case are shown in Fig. 13. We remark that close to the transition between the Mott and cluster phases ($g_{d}\approx 0.8E_{r}$) there is a strong competition between lattice localization, contact repulsion, and dipolar attraction. Furthermore, depending on the number of particles and sites used, frustration can emerge (for example 6 particles wanting to localize on two sites will not be pinned symmetrically around the central site in an optical lattice with an odd number of sites). As a consequence, the corresponding ground states can be quasidegenerate and spontaneously break the inversion symmetry of the lattice. This can lead to the states being pinned away from the central site of the optical lattice [see Fig. 12(c) and Fig. 13(b)]. Nevertheless, we have verified that pinning at different sites does not lead to any qualitative differences in the density nor in the entropy dynamics.
@@ Appendix C Initial states
Figure 12: Initial states (solid line) and optical lattice (dashed line, scaled down by 20$\times$) used in the quench protocol for $N=5$ bosons and $M=10$ orbitals in $S=5$ sites (unit filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.57E_{r}$, (c) $g_{d}=-0.82E_{r}$, (d) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
@@
Figure 13: Initial states (solid line) and optical lattice (dashed line, scaled down by 20$\times$) used in the quench protocol for $N=6$ bosons and $M=12$ orbitals in $S=3$ sites (double filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.82E_{r}$, (c) $g_{d}=-1.01E_{r}$, (d) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
Appendix D Other measures of correlation
In this appendix we present the results for the one-body and two-body Glauber correlation functions $g^{(1)}(x,x^{\prime})=\frac{\rho^{(1)}(x,x^{\prime})}{N\sqrt{\rho(x)\rho(x^{\prime})}}$ and $g^{(2)}(x,x^{\prime})=\frac{\rho^{(2)}(x,x^{\prime})}{N^{2}\rho(x)\rho(x^{\prime})}$. These quantities, shown in Figs. 14 to 17, provide a more comprehensive picture of the correlations in the system by normalizing the one-body density $\rho^{(1)}(x,x^{\prime})$ and two-body density $\rho^{(2)}(x,x^{\prime})$ with the particle number and density, although they conclude with the same physics drawn in the main text. We find that the dynamics of first-order and second-order coherence is strongly related to the many-body dynamical fragmentation. The parameters for post quench dynamics remain the same as described in the main text.
The results for the unit filling case are presented in Figs. 14 and 15. For the prequench state with $g_{d}=0.0E_{r}$ [Fig. 14(a)] and with $g_{d}=-0.45E_{r}$ [Fig. 14 (d)], the diagonal of the first order correlation function shows five completely separated coherent regions (bright lobes) where $|g^{(1)}|^{2}$ $\approx 1$. The coherence of bosons within the same well is maintained whereas the off-diagonal correlation ($x\neq x^{\prime}$) vanishes. This describes the five-fold fragmented fully localized Mott phase. The corresponding $g^{(2)}$ in Figs. 15(a) and (d) shows five completely extinguished lobes (correlation holes), i.e. the probability of finding two bosons in the same place ($x=x^{\prime})$ is zero. On the other hand, second order coherence is maintained between the wells. For $g_{d}=0.0E_{r}$, as the many body state remains five-fold fragmented throughout the post quench dynamics, the corresponding correlation dynamics also remains stable as seen in Figs. 14(b)-(c) and 15(b)-(c). For $g_{d}=-0.45E_{r}$, the initially fully localized Mott state retains stability for long time but eventually becomes fully fragmented. This behavior is exemplified by the disappearance of diagonal correlation in one-body [Fig. 14(e)-(f)] and disruption of the correlation hole in two-body [Fig. 15 (e)-(f)] coherence. However, the faded five spots across the diagonal exhibit the presence of five bosons before complete evaporation. For $g_{d}=-0.82E_{r}$, as the many-body state is clustered in two consecutive lattice sites in one side, the corresponding one-body correlation is also confined with distinct many-body features [Fig. 14 (g)]. Upon quenching, coherence is gradually lost and finally evaporates as seen in Figs. 14 (h)-(i). The corresponding two-body coherence [Fig. 15 (g)] exhibits a correlation hole confined only to the two consecutive lattice in one side and in the post quench dynamics it follows the evaporation process (Figs. 15 (h)-(i)). For $g_{d}=-2.02E_{r}$, the one-body coherence is clustered in the central lattice [Fig. 14 (j)] and the correlation hole moves to the central position as well [Fig. 15 (j)]. During the post-quench dynamics, both one- and two-body coherence are lost with clear signature of many-body features.
In Figs. 16 and 17, we plot the one- and two-body correlation functions $g^{(1)}$ and $g^{(2)}$ for the double filling case where the fermionized Mott phase is perturbed by weak attractive long-range interaction. For the prequench state with $g_{d}=0.0E_{r}$ [Fig. 16(a)], the coherence between wells is already lost as $|g^{(1)}|^{2}\approx 0$ for all off-diagonal points $(x\neq x^{\prime})$. In addition, due to double occupation, the first order coherence within each well is designed by two distinct bright lobes signifying the presence of two fermionized bosons. The second order correlation function $g^{(2)}$ has a fully developed correlation hole along the diagonal with distinct many-body features in each lattice site. During the post-quench dynamics, correlation in each well first expands and then is quickly lost [Figs. 16(b)-(c) and 17(b)-(c)]. For $g_{d}=-0.82E_{r}$, diagonal correlations are confined to two consecutive lattice sites [Fig. 16(d)]. Upon quenching, the structure of correlation is lost and only a thin thread-like diagonal correlation is retained before complete disruption (Fig 16(e)-(f)). Fig. 17(d) shows that the correlation hole stays along the diagonal of two consecutive lattice sites and then disappears with time (Fig. 17(e)-(f)) after the quench. Similarly, for $g_{d}=-2.02E_{r}$, as the atoms are centrally localized, both the diagonal one-body correlation and the correlation hole are confined in the central lattice site [Figs. 16(g) and 17(g)]. Upon quenching, correlation is lost maintaining a thin diagonal line showing the presence of bosons before total evaporation [Figs. 16(h)-(i) and 17(h)-(i)].
Figure 14: Dynamics of the Glauber one-body correlation function $g^{(1)}(x,x^{\prime})$ post quench for $N=5$ bosons and $M=10$ orbitals with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.45E_{r}$, (g)-(i) $g_{d}=-0.82E_{r}$, (j)-(l) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
Figure 15: Dynamics of the Glauber two-body correlation function $g^{(2)}(x,x^{\prime})$ post quench for $N=5$ bosons and $M=10$ orbitals with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.45E_{r}$, (g)-(i) $g_{d}=-0.82E_{r}$, (j)-(l) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
@@ Appendix D Other measures of correlation
Figure 16: Dynamics of the Glauber one-body correlation function $g^{(1)}(x,x^{\prime})$ post quench for $N=6$ bosons and $M=12$ orbitals in $S=3$ sites (double filling) with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.82E_{r}$, (g)-(i) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.

### v2

@@ VI Conclusions
We have investigated the quench dynamics of strongly-interacting bosons in an optical lattice for a wide range of dipolar interaction strengths and different fillings, generalizing the sTG quench procedure to the presence of long-range interactions and more complex initial states. Our protocol consisted in preparing an initial state of bosons with moderate contact repulsions and weaker dipolar attractions, and then quenching the strength of the contact interactions to the negative (attractive) regime. We have calculated numerous observables to study the time evolution, including one-body density, reduced one-body and two-body densities, natural occupations, and autocorrelation functions. By analyzing these quantities, we have mapped out the effect of different dipolar interaction strengths on the stability of the post-quench states.
As expected, starting from a unit-filled Mott state at zero dipolar interactions we recollect the robustness of the original sTG quench, where the localized Mott peaks remain unperturbed for a very long time. At weak dipolar interactions, this stability is also retained for comparably long times, as the Mott peaks remains localized and uncorrelated. By increasing the dipolar interactions further, however, the initial bosonic configuration progressively condenses into localized clusters. The cluster states are much more sensitive to the quench procedure and a calculation of the autocorrelation function reveals a rapid loss of fidelity after the quench. A similar fate is observed by starting from doubly filled states, be them in doubly-occupied Mott configurations (lower dipolar interactions) or cluster states and solitons (higher dipolar interactions). The state rapidly collapses under the action of the combined attractions. Overall, we found no evidence of intermediate prethermal states in the dynamics.
By examining different correlations measures, we find that the cause for this behavior lies within the structure of the initial states, that exhibit pockets of correlations within the dimers or within different peaks of the cluster states. These pockets of correlations spread upon quenching and lead to instability and eventual collapse. In particular, the comparison between single- and double-filling systems highlights the critical role of coherence in stabilizing dynamics, with the loss of off-diagonal coherences driving sensitivity to interaction strength in the single-filling case, while being suppressed in the double-filling case due to attractive dipolar interactions. This underscores the importance of coherence in determining the stability of many-body states.
@@
Our study sheds light onto the behavior and stability of strongly correlated bosonic states in the presence of long-ranged interactions. It also showcases the potential of harnessing such long-range interactions and quenching protocols to achieve novel pathways to the creation and control of excited quantum states of matter. These questions are particularly timely since the systems we have investigated should be easily engineered in state-of-the-art ultracold experiments, e.g. with magnetic quantum gases [63] or dipolar molecules that can be controlled with unprecedented precision [64]. The complex atomic structure and the large dipole moment of these systems facilitate the control of the interatomic interaction strength, specially the relative strength of contact over dipolar interactions. This will lead to further exploration of few and many-body physics.
@@
In the future, we envision that similar approaches could be employed to study the complex interplay of different energy scales introduced by the competition between localizing optical lattices, local (contact) interactions, and long-range interactions, which could lead to a multitude of exotic quantum states. Our study could also be extended to analyze the stability of other fermionization and crystallization phenomena, potentially in higher dimensions, in disordered landscapes, or fostered by other kind of long-range interactions, e.g. mediated by cavities [65, 66]. This last setup is particularly promising as it has been shown to give rise to a plethora of intriguing equilibrium and out-of-equilibrium phases of matter, such as higher-order superfluidity, Mott phases, and limit cycles [67, 68, 69, 70]. It would be interesting to analyze whether the interplay of such phases with dipolar interactions could lead to a stabilization of prethermal states.
Another interesting route to pursue is the connection with the accuracy of dipolar quantum simulators [71]. Our method solves the many-body Schrödinger equation in the continuum, and thus should be an accurate representation of the physics at play in the actual quantum simulators and not just in the simulated lattice models (e.g. extended Bose-Hubbard models) studied elsewhere in the literature. For example, a relevant question to ask is to what extend a similar interaction quench would lead the physics to deviate from that of a lattice model as a function of dipolar interaction strength, optical lattice depth, and other system parameters. It could also include the investigation of quantum many-body scar states which can give rise to infinitely long-lived coherent dynamics under quantum quenches from well controlled initial states [72].
Another open avenue for future studies is the impact of long-range interactions on quantum holonomy – where parameters are tuned to realize a cycle starting from the free system, to Tonks-Girardeau and super-Tonks-Girardeau regimes, and then back to the free system [73]. The comparison between sudden quench and adiabatic quench will help to understand the underlying correlation dynamics. Furthermore, continuous quenches could provide information on defect dynamics and criticality by means of Kibble-Zurek formalism [74, 75].
@@ Appendix A MCTDH-X
In this appendix, we briefly review the numerical method used to obtain the time evolution of the few-boson systems presented in the main text. We employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles implemented by the MCTDH-X software [51, 52, 53, 55, 54, 56]. As MCTDH-X solves the many-body Schrödinger equation in time, it is the ideal method to probe quench dynamics of interacting ultracold systems. In fact, MCTDH-X has been applied to several dipolar-interacting systems before [76, 77, 78, 79, 80, 71].
@@
Since we have access to the full many-body wave function at any time, it is a straightforward task to compute overlaps and fidelities of the type $\Psi(t^{\prime})\middle|\Psi(t)\right>$. In this work, we are in particular interested in evaluating the (absolute square of the) autocorrelation function with the initial many-body state,
$\mathcal{F}(t)\equiv|\left<\Psi(0)\middle|\Psi(t)\right>|^{2},$ (16)
which captures the stability of the initial state in time after the quantum quench, i.e. how much the time evolved state overlaps with its initial state.
@@ Appendix B System parameters
We choose the wavelengths to be compatible with real experimental realizations in ultracold atomic labs, i.e. $\lambda_{0}\approx 532.2$ nm. This gives a wave vector $k_{0}\approx 5.903\times 10^{6}$ m-1. Depending on the setup, we set hard-wall barriers to restrict the optical lattice to the center-most five minima (single-filling simulations) or three minima (double-filling simulations).
@@ B.2 Energies
The unit of energy $\bar{E}$ is effectively defined in terms of the recoil energy of the optical lattice, i.e. $E_{r}\equiv\frac{\hbar^{2}k_{0}^{2}}{2m}\approx 2.204\times 10^{-30}$ J with $m\approx 51.941$ Da the mass of 52Cr atoms, a dipolar magnetic species amply used in experiments [85, 86]. More specifically, we define the unit of energy as $\bar{E}\equiv\frac{\hbar^{2}}{m\bar{L}^{2}}=\frac{2E_{r}}{\pi^{2}}\approx 4.466\times 10^{-31}$ J.
In typical experiments, the optical lattice depth is varied in regimes of up to several tens of recoil energies (in particular when lattice physics is quantum simulated). In our simulations, we probe similar regimes of $V_{0}=100\bar{E}\approx 20E_{r}$. The on-site interactions are kept at fixed magnitude $|g_{0}|=10\bar{E}\approx 2E_{r}$, unless otherwise stated. During the quench, we change only their sign (from repulsive to attractive). When we turn on dipolar interactions, we probe regimes up to strength $|g_{d}|=2.0\bar{E}\approx 0.4E_{r}$, since the physics does not change qualitatively with stronger dipolar interactions. Empirically, we find that around $|g_{d}|\approx 0.2E_{r}$ – i.e. around 10% of the contact interaction strength – is sufficient to condense all bosons into a central droplet despite the strong initial repulsions.
@@ B.3 Time
The unit of time is also defined from the unit of length, as $\bar{t}\equiv\frac{m\hat{L}^{2}}{\hbar}=\frac{m\lambda_{0}^{2}}{\hbar}=0.5903\times 10^{-4}$ s $=59.03$ $\mu$s. In our simulations, we ran all time evolutions up until at least $t\approx 5-6\bar{t}\approx 300-350$ $\mu$s. However, for the fidelity calculations, we typically let the time evolution run for longer than $t\approx 20\bar{t}\approx 1.2$ ms.
@@
potential depth $V=100.0\bar{E}\approx 20\>E_{r}$
on-site repulsion $|g_{0}|=10.0\bar{E}\approx 2\>E_{r}$
dipolar interaction e.g. $g_{d}=-1.0\bar{E}\approx-0.2\>E_{r}$
@@
Appendix C Orbital convergence
In this section we discuss how our results depend on the number of orbitals $M$ employed in the computations. The accuracy of the MCTDH-B ansatz crucially depends on the choice of $M$. Strictly speaking, the wave function becomes exact only in the limit $M\rightarrow\infty$. However, very often it is possible to achieve converged, numerically exact results while employing a finite number of orbitals. An important criterion used to quantify the correctness of the calculated observables is orbital convergence, which requires the main features of the system at hand to stabilize as the number of orbitals used in the computations is progressively increased. To that end, we have probed convergence both in the initial state preparation and in the quench dynamics by repeating the main text calculations with $M=12$ to $M=15$ orbitals.
Let us preface the discussion of the convergence results by saying that the many-body problem we are currently investigating is extraordinarily challenging for several reasons: (1) sudden quench procedures are incredibly disruptive to the wave function structure and to the computation of time evolutions, because of the deep orbital reorganization that has to happen in real time. (2) This is especially true when we consider very strong interactions (several recoil energies) as in our case. (3) Attractive bosons are particularly difficult to converge with the MCTDH-B method. The issue is more severe with short-range interaction in free space [87], but can be mitigated for finite-range interaction [51]. (4) there is a computational tradeoff between employing a fine spatial resolution and enabling a faster integration of the many-body Schrödinger equation to reach longer times. In this work, we have chosen to focus on probing the longest times possible with the highest accuracy. This requires us to employ a coarser spatial resolution, which may sometimes result in slight asymmetries and kinks in the density profile.
We first present the density profile for the initial states used in the quench protocol and studied in the main text. The initial states for the unit filling case obtained with different numbers of orbitals can be seen in Fig. 9, while the ones for the double filling case are shown in Fig. 10. We remark that close to the transition between the Mott and cluster phases ($g_{d}\approx 0.2E_{r}$) there is a strong competition between lattice localization, contact repulsion, and dipolar attraction. As we can see, around the phase transition there are some fluctuations around the symmetric structure for lower values of $M$, see Fig. 9(b). However, employing $M=15$ yields a safely converged initial state for all the parameter ranges probed in this work.
@@ Appendix C Orbital convergence
Figure 9: Orbital convergence of the initial states (solid line) and optical lattice (dashed line, scaled down by 20$\times$) used in the quench protocol for $N=5$ bosons in $S=5$ sites (unit filling) with (a) $g_{d}=-0.1E_{r}$, (b) $g_{d}=-0.18E_{r}$, (c) $g_{d}=-0.19E_{r}$, (d) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
@@
Figure 10: Orbital convergence of the initial states (solid line) and optical lattice (dashed line, scaled down by 20$\times$) used in the quench protocol for $N=6$ bosons in $S=3$ sites (double filling) with (a) $g_{d}=-0.16E_{r}$, (b) $g_{d}=-0.17E_{r}$, (c) $g_{d}=-0.19E_{r}$, (d) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
We now discuss how the dynamics is affected by adopting an increasing number of orbitals in the MCTDH decomposition. For brevity, we only consider the single-filling case with $g_{d}=-0.18E_{r}$, as it appears to be the harder case to converge even for the initial state. We nevertheless verified that the dynamics is converged in the number of orbitals for other values of $g_{d}$ and double filling.
The density dynamics after the quench is illustrated in Fig. 11 for an increasing number of orbitals up to $M=15$ (corresponding to 3 orbitals per site on average). While we do observe some slight asymmetries with a lower number of orbitals, the results for $M=14$ and $M=15$ very closely resemble each other, indicating that the density dynamics is converged at $M=15$.
Concomitantly, we observe that the orbital occupation saturates with increasing number of orbitals, as evinced from Fig. 12. In fact, only up to eight orbitals appear to have a macroscopic population. We have verified that the additional orbitals (9-15) systematically contribute to less than 1% to the total many-body wave function throughout the entire time evolution. Since adding more orbitals not only does not change the overall dynamics, but also results in negligible occupation of the newly introduced orbitals, we can declare the overall dynamics converged.
@@
Figure 11: Post quench density dynamics for $N=5$ bosons in $S=5$ sites (single filling) with dipolar interaction strength $g_{d}=-0.18E_{r}$, plotted for increasing orbital number $M$: (a) $M=12$, (b) $M=13$, (c) $M=14$, (d) $M=15$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
Figure 12: Dynamics of orbital occupation for $N=5$ bosons in $S=5$ sites (single filling) with dipolar interaction strength $g_{d}=-0.18E_{r}$, plotted for increasing orbital number $M$: (a) $M=12$, (b) $M=13$, (c) $M=14$, (d) $M=15$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
Appendix D Autocorrelation for $M=15$
In this appendix, we show results for the calculation of the autocorrelation function with $M=15$ orbitals. The autocorrelation function is mapped as a function of the interaction strength $g_{d}$ and time $t$ in Fig. 13. The behavior of the autocorrelation function is essentially equivalent to the one obtained for $M=12$ orbitals and presented in the main text. This ensures that our results with $M=12$ orbitals are already reliable.
The autocorrelation function separates the dipolar interaction strengths into three regimes. A first regime persists down to $g_{d}\approx-0.1E_{r}$ and is characterized by exceptional long-time stability of the quenched state. A second regime can be observed between $g_{d}\approx-0.1E_{r}$ and $g_{d}\approx-0.19E_{r}$ and exhibits a reduction of stability with values of autocorrelation dropping all the way down to 0.2 for the stronger interactions. A third regime occurs below $g_{d}\approx-0.19E_{r}$ and reveals a complete collapse of the initial state as evinced by autocorrelation values close to 0.